# Mining Glitch Tokens in Large Language Models via Gradient-based Discrete Optimization

### v2

@@ Abstract
Glitch tokens in Large Language Models (LLMs) can trigger unpredictable behaviors, compromising model reliability and safety. Existing detection methods often rely on manual observation to infer the prior distribution of glitch tokens, which is inefficient and lacks adaptability across diverse model architectures. To address these limitations, we introduce GlitchMiner, a gradient-based discrete optimization framework designed for efficient glitch token detection in LLMs. GlitchMiner leverages an entropy-based loss function to quantify the uncertainty in model predictions and integrates first-order Taylor approximation with a local search strategy to effectively explore the token space. Our evaluation across various mainstream LLM architectures demonstrates that GlitchMiner surpasses existing methods in both detection precision and adaptability. In comparison to the previous state-of-the-art, GlitchMiner achieves an average improvement of 19.07% in precision@1000 for glitch token detection. By enabling efficient detection of glitch tokens, GlitchMiner provides a valuable tool for assessing and mitigating potential vulnerabilities in LLMs, contributing to their overall security. Our code is available at https://github.com/wooozihui/GlitchMiner.
@@ 1 Introduction
Large Language Models (LLMs), such as GPT-4 gpt4 , have revolutionized natural language processing, enabling breakthroughs in critical fields, including code generation LLM_code1 ; LLM_code2 ; LLM_code3 , medical research medical1 ; medical2 , and education education1 ; education2 . Despite their strengths, these models harbor a subtle yet critical vulnerability: glitch tokens geiping2024coercing . Glitch tokens are anomalous tokens that disrupt the model’s ability to understand or generate accurate responses, leading to unpredictable behaviors such as inappropriate content, repetitive errors, hallucinated outputs, or offensive language geiping2024coercing ; solidgoldmagikarp ; GlitchHunter ; Magikarp . This issue stems from the tokenization process, potentially arising due to insufficient training or poorly embedded tokens within the model’s vocabulary.
For instance, as shown in Figure 1, the Llama2-7B-chat-hf llama2 model fails to recognize the token “oreferrer,” treating it as an uncorrelated term. This oversight reveals deeper vulnerabilities in token handling, leading to unexpected outputs that can compromise the reliability of LLMs, especially in high-stakes applications. Mitigating these glitches is crucial to ensuring the trustworthiness and robustness of LLM-generated content.
Existing methods for detecting glitch tokens primarily rely on manual observation to assume their prior distribution solidgoldmagikarp ; GlitchHunter ; Magikarp . However, this approach is not only time-consuming and labor-intensive but also struggles to adapt to different model architectures and vocabularies. Moreover, it often overlooks glitch tokens that fall outside predefined distribution patterns, limiting the detection’s coverage and accuracy.
To overcome these challenges, we propose GlitchMiner, a novel gradient-based discrete optimization framework for efficiently detecting glitch tokens in LLMs. Unlike previous methods that rely on manual observation of glitch tokens, GlitchMiner focuses on identifying tokens that deviate significantly from the normal token distribution—where most tokens exhibit high certainty and predictable behavior. As illustrated in Figure 2, GlitchMiner leverages the uncertainty in token predictions to uncover potential glitches. It employs an entropy-based loss function to measure uncertainty and uses a local discrete optimization strategy, combining first-order Taylor approximation Taylor with local search, to efficiently explore the token space. This adaptive approach enables GlitchMiner to surpass the limitations of earlier methods, providing more comprehensive detection without relying on predefined distribution assumptions.
Our main contributions in this paper include:
  1. 1.
    Entropy-based Loss Function: We propose using entropy as the optimization objective, effectively capturing the uncertainty in model predictions without relying on manually defined glitch token prior distributions.
  2. 2.
@@
    Local Search Strategy: We design a novel discrete optimization method that combines first-order Taylor approximation with local search, significantly improving the efficiency and accuracy of glitch token detection.
  3. 3.
    Wide Applicability: Through experiments on multiple mainstream LLM architectures, we demonstrate the generalizability and effectiveness of GlitchMiner. Experimental results show that GlitchMiner improves the precision@1000 of glitch token detection by an average of 19.07% compared to previous state-of-the-art method.
  4. 4.
    Efficient Detection: Our method can quickly locate potential glitch tokens in large-scale vocabularies, greatly improving detection efficiency and providing a powerful tool for assessing and improving the security of LLMs.
@@
The structure of this paper is as follows: Section 2 reviews related work, including glitch token detection and gradient-based discrete optimization methods; Section 3 details the design and implementation of GlitchMiner; Section 4 presents experimental results and analysis; finally, Section 5 concludes the paper and discusses future research directions.
Figure 1: Example of a glitch token "oreferrer", causing misinterpretation in the Llama2 model.
Figure 2: Overview of the GlitchMiner Framework. This figure illustrates the two-stage process: (1) Initialization Stage – filters out unnecessary tokens to reduce the search space; (2) Mining Stage – iteratively selects batches of tokens, evaluates their entropy, and determines whether they are glitch tokens.
2 Related Work
@@ 2.1 Glitch Token Detection
Detecting glitch tokens in LLMs has garnered increasing attention. Several methods have been proposed to address this issue, including Magikarp Magikarp , GlitchHunter GlitchHunter , and GlitchProber GlitchProber , each utilizing different strategies to identify glitch tokens.
Magikarp Magikarp detects glitch tokens by analyzing token embeddings. It uses characteristic patterns, such as the $\ell_{2}$ norm, in the embedding space to identify under-trained tokens and employs simple repetition tasks to test whether the model can accurately reproduce these tokens. This method focuses on quickly detecting problematic tokens through direct queries, making it fast and lightweight.
GlitchHunter GlitchHunter adopts a clustering-based approach. It assumes that glitch tokens tend to cluster in the embedding space and constructs a Token Embedding Graph (TEG), where nodes represent tokens connected based on their embedding similarity. By applying the Leiden algorithm traag2019louvain , GlitchHunter identifies potential glitch clusters. Hypothesis testing is then used to refine these clusters, iterating multiple times to ensure the accuracy of detection.
@@
Both Magikarp and GlitchHunter rely on manually observed prior distributions to sample token space for glitch token detection, performing well when their assumptions align with the actual glitch token distribution. However, their adaptability can be limited when real-world distributions deviate from these priors. GlitchProber aims to improve detection through automation, but it requires a large number of queries to collect sufficient training data. Additionally, the need for extra training introduces computational overhead, impacting efficiency in practice.
Our method overcomes these limitations by using entropy-based loss functions to quantify uncertainty in predictions, without relying on priors or additional training.
@@ 2.2 Gradient-based Discrete Optimization
Gradient-based discrete optimization methods hotflip ; autoprompt ; GCG leverage gradient information to predict how individual tokens impact the loss function. These approaches typically treat the one-hot encoding of tokens or token embeddings as continuous vectors to compute gradients, guiding token replacements for optimization.
@@
Both AutoPrompt and GCG rely on large batch sampling to reduce inaccuracies in gradient predictions. These inaccuracies occur because Taylor approximation become less reliable when the input tokens are far from the original reference points.
To address this challenge, we introduce a local search strategy in our approach. By focusing on a smaller, localized token space, our method improves precision without the need for large batch sampling. This targeted search not only enhances precision but also avoids unnecessary queries over non-glitch tokens, significantly boosting query efficiency for glitch token detection.
@@ 3.2 Glitch Token Verification
Task Template. Previous research typically employed a repetition task to verify whether a token is a glitch token by checking if the model can repeat a given token. If the model fails to reproduce the token correctly, the token is labeled as a glitch token. However, we found that the prior task templates are not directly suitable for gradient-based discrete optimization. For instance, in Magikarp Magikarp , the token being checked is repeated multiple times, forcing the model to output the current token, which can lead to unnecessary computational overhead and imprecise gradient calculations of the input token.
To address this issue, we designed a more straightforward task template as follows:
@@
In this template, the blue text represents the prompt used to verify glitch tokens. The green text, {token}, corresponds to the input token being tested, while the red text, {token}, represents the model’s predicted output. If the predicted token does not match the input token, the input token is classified as a glitch token.
The response from the assistant, starting with “Sure, the string is:"«”, is prefilled as part of the template, ensuring consistency in the model’s response format. The token is wrapped in the "«" and "»" symbols, similar to the token filter, to prevent interference from other characters during token encoding.
@@ 3.3 Token Selection
Entropy as Loss Function. One of the key distinctions between our approach and previous work lies in our token selection strategy. Prior methods often rely on manually observed prior knowledge about the distribution of glitch tokens to select those most likely to be glitch tokens. However, this approach can overlook tokens that do not fit the prior assumptions but are still glitch tokens. Ultimately, the effectiveness of these methods depends on how closely the manually defined priors align with the actual distribution of glitch tokens.
Given that manual priors of glitch token distribution can be inaccurate, a more robust approach is to identify tokens that deviate significantly from the normal token distribution. To achieve this, we propose using entropy as the loss function in our optimization process. Entropy measures the uncertainty or randomness in a probability distribution. For normal tokens, we expect the model to predict the next token with high confidence, resulting in a low-entropy distribution. Maximizing entropy helps us find tokens that cause more uncertainty in the model’s predictions, thereby identifying tokens that deviate from the expected behavior of normal tokens. This allows us to discover glitch tokens without depending on observed distribution patterns or manually defined assumptions about their characteristics.
Optimization Objective. Based on the above ideas, we present the mathematical description of the optimization objective. In each iteration, the goal is to find a batch of tokens with the highest entropy from the set of unverified tokens. Let $\mathcal{T}_{c}$ represent the current candidate set, defined as $\mathcal{T}_{c}=\mathcal{T}\setminus(\mathcal{T}^{*}\cup\mathcal{G})$.
Let $\mathbf{h}(t)$ represent the context embedding for a token $t$, and let $P(v\mid\mathbf{h}(t))$ denote the probability distribution over the vocabulary $\mathcal{V}$, predicted by the model for token $t$. The entropy $H(t)$ for token $t$ is defined as:
@@
In each iteration, the optimization objective is to find a batch of tokens $\mathcal{B}$ that maximizes the entropy $H(t)$, i.e.,
@@
Local Search Strategy. To efficiently solve the aforementioned optimization problem, we propose a local search strategy based on AutoPrompt. This approach combines local search with first-order Taylor approximation to estimate the entropy of candidate tokens, enabling us to select the batch most likely to contain glitch tokens.
Our strategy initiates by selecting an initial token $t_{0}$, computing its entropy $H(t_{0})$, and the corresponding gradient $\nabla_{e}H(t_{0})$, where $\nabla_{e}H(t_{0})$ represents the partial derivative with respect to the embedding vector $e_{t_{0}}$. For each candidate token $t\in\mathcal{T}_{c}$, we estimate its entropy using the first-order Taylor approximation:
@@
where $e_{t}$ and $e_{t_{0}}$ denote the embedding vectors of tokens $t$ and $t_{0}$, respectively. This approximation estimates the change in entropy for the candidate token $t$ based on the change in its embedding relative to $e_{t_{0}}$, guiding the selection of tokens with potentially higher entropy.
To improve the accuracy of the approximation, we introduce a local search strategy. We define $\mathcal{N}_{K}(t_{0})$ as the set of $K$ nearest neighbor tokens to $t_{0}$:
$\mathcal{N}_{K}(t_{0})=\{t\in\mathcal{T}_{c}:t\text{ is one of the }K\text{ nearest neighbors of }t_{0}\}$
Within this local neighborhood $\mathcal{N}_{K}(t_{0})$, we select the batch $\mathcal{B}$ of $B$ tokens with the highest approximated entropy:
$\mathcal{B}=\arg\max_{\mathcal{B}\subset\mathcal{N}_{K}(t_{0}),|\mathcal{B}|=B}\sum_{t\in\mathcal{B}}\hat{H}(t)$
After selecting the batch $\mathcal{B}$, we compute the actual entropy values $H(t)$ for each token in the batch and choose the token with the highest actual entropy as the starting point $t_{0}$ for the next iteration. This ensures that each iteration is guided by the token with the most promising entropy, allowing the optimization process to converge more efficiently.
This local optimization strategy effectively mitigates the issue of inaccuracies in the Taylor approximation when applied globally while maintaining computational efficiency. By focusing on a local neighborhood in each iteration, we can more accurately estimate the entropy of candidate tokens, thereby enhancing the efficiency of glitch token detection. This approach strikes a balance between exploration of the token space and exploitation of local information, facilitating a more targeted and effective search for glitch tokens.

### v3

@@ Abstract
Glitch tokens in Large Language Models (LLMs) can trigger unpredictable behaviors, threatening model reliability and safety. Existing detection methods often depend on predefined patterns, limiting their adaptability across diverse LLM architectures. We propose GlitchMiner, a gradient-based discrete optimization framework that efficiently identifies glitch tokens by leveraging entropy to quantify prediction uncertainty and a local search strategy for exploring the token space. Experiments across multiple LLM architectures show that GlitchMiner outperforms existing methods in both detection accuracy and adaptability, achieving over 10% average efficiency improvement. GlitchMiner enhances vulnerability assessment in LLMs, contributing to more robust and reliable applications. Code is available at https://github.com/wooozihui/GlitchMiner.
@@ 1 Introduction
Large Language Models (LLMs) have revolutionized natural language processing, enabling breakthrough applications in fields ranging from code generation LLM_code1 ; LLM_code2 ; LLM_code3 to medical research medical1 ; medical2 and education education1 ; education2 . However, the increasing reliance on these models in critical domains has brought to light a subtle yet significant vulnerability: glitch tokens geiping2024coercing . These anomalous tokens can severely disrupt an LLM’s ability to generate coherent and reliable outputs, potentially leading to repetitive errors, nonsensical responses, or even harmful content. The implications of such disruptions are particularly concerning in high-stakes applications where model reliability is paramount.
Recent research efforts have made strides in understanding and detecting glitch tokens. Approaches like Magikarp Magikarp and GlitchHunter GlitchHunter have provided valuable insights, focusing on abnormal embedding patterns and clustering properties of these tokens. However, these methods are inherently limited by their reliance on manually observed patterns and predefined distribution assumptions. This dependence not only constrains their adaptability across diverse LLM architectures but also potentially overlooks novel forms of glitch tokens that do not conform to expected patterns.
In this paper, we introduce GlitchMiner, a novel framework that fundamentally reimagines the approach to glitch token detection. Moving beyond pattern-based heuristics, GlitchMiner employs gradient-based discrete optimization to directly analyze an LLM’s prediction behavior. At the core of our method is the use of entropy as an exploratory metric, guiding our search through the vast token space to identify regions of unexpected model behavior.
In summary, the key innovations of GlitchMiner include:
  • Entropy-Based Optimization: We introduce entropy as a novel optimization objective for glitch token detection. This approach allows us to systematically identify tokens that induce high uncertainty in the model’s predictions, moving beyond predefined patterns or embedding anomalies. By focusing on the model’s prediction behavior, we can detect a broader spectrum of glitch tokens, including those that might be missed by traditional embedding-based methods.
@@
  • Gradient-Guided Local Search: Our method combines first-order Taylor approximation with a local search strategy, efficiently navigating the vast token space. This approach significantly improves both the efficiency and accuracy of glitch token detection, allowing us to explore the token space more thoroughly while maintaining computational feasibility.
  • Adaptive Cross-Architecture Detection: GlitchMiner’s framework is designed to adapt to the unique characteristics of different LLM architectures without relying on model-specific assumptions. This flexibility enables effective glitch token detection across a wide range of models, addressing a key limitation of previous methods that often relied on architecture-specific heuristics.
We conducted extensive experiments across 10 popular LLM architectures, covering models from the Llama, Qwen, Gemma, Phi, and Mistral families. The results show that GlitchMiner surpasses existing methods in 90% of cases when it comes to glitch token detection efficiency. Specifically, GlitchMiner achieves an average Detected@1000 score of 612.0 and a Detected@2000 score of 980.1. These results represent improvements of 10.5% and 10.7% respectively over the previous leading methods. In the context of glitch token detection, these improvements are significant as they translate to hundreds of additional potentially harmful tokens being identified. Such enhancements not only improve detection performance but also optimize query efficiency, which is crucial for enhancing the reliability and robustness of LLMs in critical applications where even a single undetected glitch token could lead to severe consequences.
@@
The remainder of this paper is structured as follows: Section 2 provides a comprehensive review of related work in glitch token detection and gradient-based optimization techniques. Section 3 details the design principles and implementation of GlitchMiner. Section 4 presents our experimental results and analysis, demonstrating the effectiveness of our approach across various LLM architectures. Finally, Section 5 concludes with a discussion of the implications of our work and outlines promising directions for future research in this critical area of LLM security and reliability.
2 Background and Related Work
2.1 What is a Glitch Token?
Glitch tokens, a concept that gained attention following the discovery of the "SolidGoldMagikarp" phenomenon by Rumbelow et al. solidgoldmagikarp , are anomalous tokens in LLMs that can trigger unexpected and often erroneous behaviors when processed by the model. These tokens typically result from irregularities in the training process, such as underrepresentation in the training data or inconsistencies in tokenization Magikarp ; geiping2024coercing . Glitch tokens have been found to exhibit several key characteristics:
  • They often have abnormal embedding patterns, such as unusually large L2 norms or atypical positions in the embedding space Magikarp .
  • When input to an LLM, they can cause the model to produce repetitive, nonsensical, or completely unrelated outputs solidgoldmagikarp ; GlitchHunter .
  • They may lead to significant deviations in the model’s internal activations compared to normal tokens GlitchProber .
  • Glitch tokens can sometimes bypass content filters or trigger unexpected model behaviors, potentially compromising the safety and reliability of LLMs geiping2024coercing .
The study of glitch tokens is crucial for understanding and mitigating potential vulnerabilities in LLMs, as these tokens can potentially be exploited for adversarial attacks or lead to unintended model behaviors in critical applications geiping2024coercing .
@@ 2.2 Glitch Token Detection
The detection of glitch tokens in LLMs has become an increasingly important area of research, given their potential to disrupt model performance and reliability. The fundamental approach to identifying these anomalous tokens often relies on repetition tasks, a method popularized by the SolidGoldMagikarp study solidgoldmagikarp . In these tasks, the model is prompted to repeat a given token, with the assumption that normal tokens should be easily reproducible, while glitch tokens often lead to failures in this simple task.
Building upon this basic principle, researchers have developed several sophisticated methods to detect glitch tokens more efficiently and accurately:
Magikarp solidgoldmagikarp pioneered a fast and lightweight approach by analyzing token embeddings. It identifies potentially problematic tokens by examining characteristic patterns in the embedding space, such as unusual $\ell_{2}$ norms. These candidate tokens are then verified using repetition tasks, allowing for quick detection through direct queries to the model.
GlitchHunter GlitchHunter introduced a clustering-based method. This approach is founded on the observation that glitch tokens often cluster together in the embedding space. GlitchHunter constructs a Token Embedding Graph (TEG) to represent token relationships and applies the Leiden algorithm traag2019louvain to identify potential glitch clusters. These clusters are then refined through iterative hypothesis testing to improve detection accuracy.
@@
While these methods have significantly advanced our ability to detect glitch tokens, they are inherently limited by their reliance on pattern-based heuristics. Magikarp and GlitchHunter focus on specific embedding patterns or clustering behaviors, while GlitchProber examines predefined patterns in activation spaces. This dependence on pattern-based approaches can constrain their adaptability across diverse LLM architectures and potentially overlook novel forms of glitch tokens.
Our proposed method, GlitchMiner, aims to address these limitations by employing gradient-based discrete optimization to directly analyze an LLM’s prediction behavior, potentially offering a more flexible and comprehensive approach to glitch token detection.
@@ 2.3 Gradient-based Discrete Optimization
Gradient-based discrete optimization methods hotflip ; autoprompt ; GCG ; PEZ leverage gradient information to predict how individual tokens impact the loss function. These approaches typically treat the one-hot encoding of tokens or token embeddings as continuous vectors to compute gradients, guiding token replacements for optimization.
@@
AutoPrompt and GCG both rely on large batch sampling to mitigate inaccuracies in gradient prediction. We identified that these inaccuracies arise from the inaccuracy of Taylor expansions when input tokens are distant from the original points. This overlooks a fundamental condition of Taylor approximation: its accuracy is highest for points close to the reference point.
Building on these works, we introduce a local search strategy in our approach. This improvement enables us to achieve high precision in gradient estimation without relying on large batch sampling, by focusing on a smaller, localized token space. By addressing the core issue of Taylor approximation accuracy, our method allows for more efficient and accurate exploration of the token space, which is particularly valuable for glitch token detection.
@@ 3.2 Glitch Token Verification
Our glitch token verification process consists of two stages: optimization-driven detection and cross-verification.
Template for Optimization. For our primary detection and optimization process, we designed the following task template:
@@
In this template, the blue text represents the prompt, the green text {token} is the input token being tested, and the red text {token} represents the model’s predicted output. If the predicted token does not match the input token, it is initially classified as a potential glitch token.
This template is specifically designed to support our gradient-based optimization process. Its simplicity allows for clear gradient computations, while the prefilled assistant response ensures a uniform starting point for each token evaluation. Wrapping the token in "«" and "»" symbols minimizes interference from surrounding context. These features enable precise entropy calculations and gradient estimations, which are crucial for our entropy-guided search strategy described in the token selection section.
Cross-Verification. While our primary template is optimized for the search process, relying solely on one template may lead to false positives. To enhance the robustness of our detection, we implement a cross-verification step using two additional templates derived from GlitchHunter and Magikarp. Each potential glitch token identified by our primary template undergoes verification with these additional templates. A token is confirmed as a glitch token only if it fails the repetition task across all templates.
This two-stage approach combines the efficiency of our optimization-driven search with the reliability of multi-template verification, significantly reducing false positives while maintaining the effectiveness of our gradient-based detection method.
@@ 3.3 Token Selection
Motivation for Entropy-Based Approach. Previous methods for glitch token detection often rely on manually observed patterns or predefined distribution assumptions. However, these approaches can be limited in their ability to adapt to diverse LLM architectures and may overlook novel forms of glitch tokens. To address these limitations, we propose an entropy-based approach. Entropy, as a measure of uncertainty in probability distributions, offers a model-agnostic way to identify tokens that cause unexpected behavior in LLMs. By focusing on tokens that maximize entropy, we can detect glitch tokens that deviate significantly from normal token behavior, regardless of their specific characteristics or the underlying model architecture.
Entropy-Guided Exploration. We define the entropy $H(t)$ for a token $t$ as:
@@
where $\mathbf{h}(t)$ is the context embedding for token $t$, and $P(v\mid\mathbf{h}(t))$ is the model’s predicted probability distribution over the vocabulary $\mathcal{V}$. By maximizing entropy, we aim to find tokens that cause high uncertainty in the model’s predictions, potentially indicating glitch behavior.
Optimization Objective. Our goal is to find a batch of tokens $\mathcal{B}$ that maximizes the total entropy:
@@
Here, $\mathcal{T}_{c}$ represents the current candidate set, defined as $\mathcal{T}_{c}=\mathcal{T}\setminus(\mathcal{T}^{*}\cup\mathcal{G})$, where:
  • $\mathcal{T}$ is the initial set of all tokens in the model’s vocabulary.
  • $\mathcal{T}^{*}$ is the set of tokens that have been filtered out or verified as non-glitch tokens. This includes special tokens, undecoded tokens, and tokens that have been checked and confirmed as normal.
  • $\mathcal{G}$ is the set of identified glitch tokens discovered so far in the process.
By excluding $\mathcal{T}^{*}$ and $\mathcal{G}$ from $\mathcal{T}$, we ensure that our search focuses only on the remaining unverified tokens, improving efficiency and preventing redundant checks.
Local Search Strategy. To efficiently solve this optimization problem while addressing the limitations of global Taylor approximation, we introduce a local search strategy. We begin by selecting an initial token $t_{0}$ and computing its entropy $H(t_{0})$ and gradient $\nabla_{e}H(t_{0})$. We then define a local neighborhood $\mathcal{N}_{K}(t_{0})$ consisting of the $K$ nearest neighbors to $t_{0}$ in the embedding space. This localized approach allows us to refine the search within a region close to $t_{0}$, improving approximation accuracy and ensuring efficient exploration of potential glitch tokens.
Within this local neighborhood, we estimate the entropy of each candidate token $t\in\mathcal{N}_{K}(t_{0})$ using first-order Taylor approximation:
@@
where $e_{t}$ and $e_{t_{0}}$ are the embedding vectors of tokens $t$ and $t_{0}$, respectively. This approximation allows us to efficiently estimate the entropy change for nearby tokens without the computational cost of exact calculations.
Based on these estimates, we select a batch $\mathcal{B}$ of $B$ tokens with the highest approximated entropy. For this selected batch, we then compute the actual entropy values and choose the token with the highest entropy as the starting point for the next iteration. This iterative process ensures that our search is guided by the most promising candidates at each step.
This local search strategy significantly improves the accuracy of entropy estimation by focusing on tokens close to the reference point, addressing a key limitation of previous gradient-based methods. By balancing exploration of the token space with exploitation of local information, our approach enables more efficient and effective glitch token detection.
The combination of entropy-guided exploration and local search allows our method to adapt to diverse LLM architectures without relying on model-specific assumptions. This results in a more robust and versatile approach to glitch token detection, capable of identifying a wide range of anomalous tokens across different model types.